OpenStudy (anonymous):

Solve for r: |5r - 18| = 0. A. 3.6 B. {3.6, -3.6} C. No Solution

OpenStudy (anonymous):

For |5r - 18| = 0. 5r - 18 =0 or -(5r -18 ) =0 5r = 18 or -5r = -18 r = 3.6 or r = 3.6 I think you can choose the suitable choice now.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome.